OpenStudy (anonymous):

volume of 10 drops of a liquid is 0.1 fluid ounce. What is the volume of 100 drops

rvc (rvc):

i think its 1 fluid ounce. maybe..

OpenStudy (anonymous):

10/0.1=100/x x=1 fluid ounce